yearly production of a 4 foot orange tree is 34 pounds of oranges. A 15 foot tree produces 75 pounds. Let h represent the height of an orange
tree and p the number of pounds of oranges produced. use formula to answer questions.
formula is m=p2-p1/h2-h1
a. list ordered pair(h.p) first ordered pair of first tree?

Answer by stanbon(75887)

yearly production of a 4 foot orange tree is 34 pounds of oranges. A 15 foot tree produces 75 pounds. Let h represent the height of an orange
tree and p the number of pounds of oranges produced. use formula to answer questions.
formula is m=p2-p1/h2-h1
m = (75-34)/(15-4) = 41/11
----------------
a. list ordered pair(h.p) first ordered pair of first tree?
Ans: (4,34)
================
Cheers,
Stan H.
================